# Model Tensor Planning

### v1

@@ 1 Introduction
Evolutionary Strategies (ES) (Hansen, 2016; Wierstra et al., 2014; Salimans et al., 2017) have also been applied in sampling-based MPC settings to improve sampling exploration. While they improve over purely local strategies in some tasks, our experiments (cf. Section 4.1) reveal that ES still fails to systematically explore multimodal control landscapes, often yielding inconsistent performance on tasks requiring long-term coordinated actions.
@@ 2 Preliminary
Notations and Assumptions. We consider the problem of sampling-based MPC. Given a dynamics model $\dot{{\bm{x}}}=f({\bm{x}},{\bm{u}})$, we consider the path sampling problem in the control space ${\mathbb{U}}\subset{\mathbb{R}}^{n}$ having $n$-dimensions at the current system state ${\bm{x}}\in{\mathbb{X}}\subset{\mathbb{R}}^{d}$. Typically, a batch of control trajectories is sampled, rolled out through the dynamics model, and evaluated using a cost function. Let a control path be $u:[0,1]\to{\bm{u}},\,{\bm{u}}(t)\in{\mathbb{U}}$, we can define the path arc length as
@@
Figure 1: Comparison of MTP interpolation methods versus CEM on PushT environment. The cost of PushT is the sum of the position and orientation error to the green target (without the guiding contact cost), and the initial T pose is randomized. The first row depicts the cost convergence over $10$ seeds. In most seeds, CEM struggles to push the object due to the lack of exploration (e.g., mode collapsing), while MTP variants always find the correct contact point to achieve the task. Note that the control magnitude of MTP is high due to the global explorative samples (see second & third rows), compared to the white noise samples (blue). B-spline helps regulate the control magnitude due to its barycentric weightings, while retaining exploration behaviors. The last row illustrates the control trajectories between $64$ tensor samples and $64$ white noise trajectories. Experiment videos and open-source implementations are publicly available at https://sites.google.com/view/tensor-sampling/ and https://github.com/anindex/mtp.
@@ 2.1 Cross-Entropy Method for Sampling-based MPC
Consider a discretized dynamical system ${\bm{x}}_{t+1}=f({\bm{x}}_{t},{\bm{u}}_{t}),\quad t=0,\dots,T-1$ with horizon $T$, where ${\bm{x}}_{t}\in{\mathbb{R}}^{d},{\bm{u}}_{t}\in{\mathbb{R}}^{n}$ denotes the state the control at time step $t$. The objective is to minimize a cumulative cost function
@@
where ${\bm{\tau}}=[{\bm{x}}_{0},\ldots,{\bm{x}}_{T}]\in{\mathbb{R}}^{(T+1)\times d},{\bm{U}}=[{\bm{u}}_{0},\ldots,{\bm{u}}_{T-1}]\in{\mathbb{R}}^{T\times n}$ are the dynamics rollout, $c({\bm{x}}_{t},{\bm{u}}_{t})$ is the immediate cost at each time step, and $c_{T}({\bm{x}}_{T})$ is the terminal cost.
@@
Elite Selection. Choose the top-$E\leq B$ elite candidates of control sequences that have the lowest cost, forming an elite set =$\mathcal{E}=\{{\bm{U}}^{(k)}\}_{k=1}^{E}$.
@@
${\bm{\mu}}_{\textrm{new}}=\frac{1}{B}\sum_{{\bm{U}}\in\mathcal{E}}{\bm{U}},\quad{\bm{\sigma}}_{\textrm{new}}^{2}=\frac{1}{B-1}\sum_{{\bm{U}}\in\mathcal{E}}({\bm{U}}-{\bm{\mu}}_{\textrm{new}})^{2}.$ (4)
@@ 2.2 Spline-based Controls
Splines provide a powerful representation for trajectory generation in MPC due to their flexibility, continuity properties, and ease of parameterization (Jankowski et al., 2023; Carvalho et al., 2024). In this work, we focus on spline-parametrization of control trajectories. Spline-based trajectories ensure smooth and feasible control inputs that satisfy constraints and objectives inherent to MPC frameworks.
@@
  • Knots. A knot $t_{i}\in[0,T]$ is a time point where polynomial pieces join. We have a non-decreasing sequence of knots $0=t_{1}\leq\ldots\leq t_{M}=T$, which partition the time interval $[0,T]$ into pieces $[t_{i},t_{i+1}]$ so that the path is polynomial in each piece. We may often have double or triple knots, meaning that several consecutive knots $t_{i}=t_{i+1}$ are equal, especially at the beginning and end, as this can ensure boundary conditions for zero higher-order derivatives.
@@
where $B_{i,p}:{\mathbb{R}}\to{\mathbb{R}}$ maps the time $t$ to the weighting of the $i$th control point, depicting the $i$th control point weight for blending (i.e., as with a probability distribution over $i$). Hence, the path point ${\bm{u}}(t)$ is always in the convex hull of control points. The B-spline functions $B_{i,p}(t)$ are fully specified by a non-decreasing series of time knots $0=t_{1}\leq\ldots\leq t_{M}=T$ and the integer polynomial degree $p\in\{0,1,\ldots\}$ by
@@
$B_{i,0}$ are binary indicators of $t\in[t_{i},t_{i+1}]$ with $11\leq i\leq M-$ 1. The 1st-degree B-spline functions $B_{i,1}$ have support in $t\in[t_{i},t_{i+2}]$ with $21\leq i\leq M-$ 2, such that $\sum_{i=1}^{M-2}B_{i,1}(t)=1$ holds. In general, degree $p$ B-spline functions $B_{i,p}$ have support in $t\in[t_{i},t_{i+p+1}]$ with $11\leq i\leq M-p-$ 1. We need $K=M-p-1$ control points ${\bm{z}}_{1:K}$, which ensures the normalization property $\sum_{i=1}^{K}B_{i,p}(t)=1$ for every degree.
@@
Motivation. Spline representation provides several benefits. (i) It ensures smooth control trajectories with guaranteed continuity of positions, velocities, and accelerations under mild assumptions of the dynamics. (ii) Spline simplifies complex trajectories through a few control points and efficiently incorporates constraints and boundary conditions. (iii) It enables easy numerical optimization thanks to differentiable and convex representations.
@@ 3.1 Tensor Sampling
Figure 2: Illustration of different tensor interpolations on evenly spaced graph with $M=3,N=9$. With a higher B-spline degree, the control trajectories exhibit more smoothness and conservative behavior, while Akima-spline aggressively and smoothly tracks control-waypoints. Note that we do not consider boundary conditions for B-spline interpolation in tensor sampling.
Inspired by (Le et al., 2024), we utilize the random multipartite graph as a tensor discretization structure to approximate global path sampling.
@@
Consider a graph ${\mathcal{G}}(M,N)=(V,E)$ on control space ${\mathbb{U}}$, the node set $V=\{L_{i}\}_{i=1}^{M}$ is a set of $M$ layers. Each layer $L_{i}=\{{\bm{u}}_{j}\in{\mathbb{U}}\mid{\bm{u}}_{j}\sim U({\mathbb{U}})\}_{j=1}^{N}$ contains $N$ control-waypoints sampled from a uniform distribution over control space (i.e., bounded by control limits). The edge set $E$ is defined by the union of (forward) pair-wise connections between layers
@@
Sampling from ${\mathcal{G}}(M,N)$. The graph nodes are represented as the control-waypoint tensor for all layers ${\bm{Z}}\in{\mathbb{R}}^{M\times N\times n}$, within the control limits. To sample a batch of $B\in{\mathbb{N}}^{+}$ control paths with a horizon $T$, we subsample with replacement ${\bm{C}}\in{\mathbb{R}}^{B\times M\times n}$ from the set of all combinatorial paths in ${\mathcal{G}}$ (cf. Algorithm 1) and further interpolate ${\bm{C}}$ into control trajectories ${\bm{U}}\in{\mathbb{R}}^{B\times T\times n}$ with different smooth structures, e.g., using Eq. 6 or Eq. 8. Sampling with replacement is cheap ${\mathcal{O}}(MN)$, while sampling without replacement typically takes ${\mathcal{O}}(N^{M})$. Since we have $N^{M}$ combinatorial paths in ${\mathcal{G}}(M,N)$, each path in ${\mathcal{G}}(M,N)$ has uniform $1/N^{M}$ mass, and the probability of sampling the same paths is small, which does not harm diversity in practice (see last row in Fig. 1).
@@
${\bm{I}}\leftarrow\texttt{randint}((B,M),1,N)$. // batch sample with replacement from ${1,\ldots,N}$ with shape $(B,M)$
@@
Control Path Interpolation. Straight-line interpolation can be realized straightforwardly by simply probing an $H=\lfloor T/M\rfloor$ number of equidistant points between layers, forming the linear coverage trajectories ${\bm{U}}\in{\mathbb{R}}^{B\times T\times n}$. However, there exist discontinuities at control waypoints using linear interpolation. Hence, we motivate spline tensor interpolations for sampling smooth control paths (cf. Fig. 2).
@@
Given two time sequences $t_{i}=i/(M+p+1),\,i\in\{0,\ldots,M+p\}$ and $t_{j}=j/T,\,j\in\{0,\ldots,T-1\}$, the B-Spline matrix ${\bm{B}}_{p}\in{\mathbb{R}}^{M\times T}$ can be constructed recursive following Section 2.2, with index $i,j$ corresponding to the element ${\bm{B}}_{i,p}(t_{j})$. Then, the control trajectories can be interpolated by performing einsum on the $M$ dimension of ${\bm{B}}_{p}\in{\mathbb{R}}^{M\times T},{\bm{C}}\in{\mathbb{R}}^{B\times M\times n}$, resulting ${\bm{U}}\in{\mathbb{R}}^{B\times T\times n}$.
@@
Definition 3 (Akima-spline Control Trajectories (Le et al., 2024)).
Given the time sequences $t_{i}=i/M,\,i\in\{0,\ldots,M-1\}$ representing the $M$ layer time slices and the control points ${\bm{C}}\in{\mathbb{R}}^{B\times M\times n}$, the Akima polynomial parameters ${\bm{A}}\in{\mathbb{R}}^{B\times(M-1)\times 4\times n}$ can be computed following Eq. 10 in batch. Then, given the time sequence $t_{j}=j/T,\,j\in\{0,\ldots,T-1\}$, the control trajectories ${\bm{U}}\in{\mathbb{R}}^{B\times T\times n}$ are interpolated following polynomial interpolation Eq. 8.
@@ 3.2 Path Coverage Guarantee
We analyze the path coverage property of ${\mathcal{G}}(M,N)$ for sampling control paths with linear interpolation. In particular, we investigate that any feasible path in the control space can be approximated arbitrarily well by a path in the random multipartite graph ${\mathcal{G}}(M,N)$ as the number of layers $M$ and the number of waypoints per layer $N$ approaches infinity.
@@
Let $u\in{\mathbb{F}}$ be any control path and ${\mathcal{G}}(M,N)$ be a random multipartite graph with $M$ layers and $N$ uniform samples per layer (cf. Definition 1). Assuming a time sequence (i.e., knots) $10=t_{1}<t_{2}<\ldots<t_{M}=$ 1 with equal intervals, associating with layers $L_{1},\ldots,L_{M}\in{\mathcal{G}}(M,N)$ respectively, then
@@
In intuition, Theorem 1 states the support of path distribution ${\mathcal{G}}(M,N)$ approximates ${\mathbb{F}}$ and converges to ${\mathbb{F}}$ as $M,N\to\infty$. Thus, sampling paths from ${\mathcal{G}}(M,N)$ provides a tensorized mechanism to efficiently sample any possible path from ${\mathbb{F}}$, which allows vectorized sampling.
@@
To quantify the exploration level of tensor sampling, one standard way is to investigate its path distribution entropy. In intuition, when $M,N\to\infty$, tensor sampling entropy also approaches infinity due to uniform sampling per layer, which is further discussed in Section A.2.
@@ 3.3 Algorithm
Here, we present the overall algorithm combining tensor and local sampling with smooth structure options in Algorithm 2. We propose a simple mixing mechanism with $\beta\in[0,1]$ by concatenating explorative and exploitative samples, forming a control trajectory tensor ${\bm{U}}$ (Line 4-8). We include the current nominal control for system stability at the fixed-point states (e.g., for tracking tasks) (Howell et al., 2022). Using simulators that allow for parallel runs (Todorov et al., 2012; Makoviychuk et al., 2021), rollout and cost evaluation can be efficiently vectorized (Line 9).
To tame the noise induced by tensor sampling, we modify the CEM update with softmax weighting on the elite set, for computing the new weighted control means and covariances similar to the MPPI update rule (Williams et al., 2017). We observe that this greatly smoothens the update over timesteps (Line 11-13) (cf. Fig. 9). Finally, similar to Howell et al. (2022), we send the first control of the best candidate, since this control trajectory is evaluated in the simulator rather than the updated mean ${\bm{\mu}}$. Notice that we have fixed tensor shapes based on hyperparameters, for all subroutines of sampling, rolling out, and control distribution updates. Algorithm 2 can be JAX jit and vmap over a number of $R$ model perturbations $\{f_{j}({\bm{x}},{\bm{u}})\}_{j=1}^{R}$, for efficient online domain randomization, while maintaining real-time control (cf. Section A.5).
@@
1 Choose interpolation type Linear, or B-spline (Definition 2), or Akima (Definition 3).
@@
5       Sample control waypoints ${\bm{C}}\in{\mathbb{R}}^{P\times M\times n}$ with $P=\lfloor\beta B\rfloor$, using Algorithm 1.
@@
10       Sum cost ${\bm{s}}=\sum_{t}{\bm{S}}_{:,t}\in{\mathbb{R}}^{B}$ and sort top-$E$ elite candidate indices ${\bm{i}}_{E}$.

### v2

@@ 1 Introduction
Evolutionary Strategies (ES) (Hansen, 2016; Wierstra et al., 2014; Salimans et al., 2017) have also been applied in sampling-based MPC settings to improve sampling exploration. While they improve over purely local strategies in some tasks, our experiments (cf. Section˜4.1) reveal that ES still fails to systematically explore multimodal control landscapes, often yielding inconsistent performance on tasks requiring long-term coordinated actions.
@@ 2 Preliminary
Notations and Assumptions. We consider the problem of sampling-based MPC. Given a dynamics model $\dot{{\bm{x}}}=f({\bm{x}},{\bm{u}})$, we consider the path sampling problem in the control space ${\mathbb{U}}\subset{\mathbb{R}}^{n}$ with the control ${\bm{u}}\in{\mathbb{U}}$ having $n$-dimensions at the current system state ${\bm{x}}\in{\mathbb{X}}\subset{\mathbb{R}}^{d}$. Typically, a batch of control trajectories is sampled, rolled out through the dynamics model, and evaluated using a cost function. Let a control path be $u:[0,1]\to{\bm{u}},\,{\bm{u}}(t)\in{\mathbb{U}}$, we can define the path arc length as
@@
Figure 1: Comparison of MTP interpolation methods versus CEM on PushT environment. The cost of PushT is the sum of the position and orientation error to the green target (without the guiding contact cost), and the initial T pose is randomized. The first row depicts the cost convergence over 101010 seeds. In most seeds, CEM struggles to push the object due to the lack of exploration (e.g., mode collapsing), while MTP variants always find the correct contact point to achieve the task. Note that the control magnitude of MTP is high due to the global explorative samples (see second & third rows), compared to the white noise samples (blue). B-spline helps regulate the control magnitude due to its barycentric weightings, while retaining exploration behaviors. The last row illustrates the control trajectories between 646464 tensor samples and 646464 white noise trajectories. Experiment videos are publicly available at https://sites.google.com/view/tensor-sampling/.
@@ 2.1 Cross-Entropy Method for Sampling-based MPC
Consider a discretized dynamical system ${\bm{x}}_{t+1}=f({\bm{x}}_{t},{\bm{u}}_{t}),\,t=0,\dots,T-1$ with horizon $T$, where ${\bm{x}}_{t}\in{\mathbb{R}}^{d},{\bm{u}}_{t}\in{\mathbb{R}}^{n}$ denotes the state the control at time step $t$. Given the state cost function $c({\bm{x}},{\bm{u}})$ and the terminal cost $c_{T}({\bm{x}})$, the objective is to minimize a cumulative cost function
@@
where the state trajectory ${\bm{\tau}}=[{\bm{x}}_{0},\ldots,{\bm{x}}_{T}]\in{\mathbb{R}}^{(T+1)\times d}$, and control trajectory ${\bm{U}}=[{\bm{u}}_{0},\ldots,{\bm{u}}_{T-1}]\in{\mathbb{R}}^{T\times n}$ are the dynamics rollout, $c({\bm{x}}_{t},{\bm{u}}_{t})$ is the immediate cost at each time step, and $c_{T}({\bm{x}}_{T})$ is the terminal cost.
@@
[…]
@@
${\bm{\mu}}_{\textrm{new}}=\frac{1}{E}\sum_{{\bm{U}}\in\mathcal{E}}{\bm{U}},\quad{\bm{\sigma}}_{\textrm{new}}^{2}=\frac{1}{E-1}\sum_{{\bm{U}}\in\mathcal{E}}({\bm{U}}-{\bm{\mu}}_{\textrm{new}})^{2}.$ (4)
@@ 2.2 Spline-based Controls
Splines provide a powerful representation for trajectory generation in MPC due to their flexibility, continuity properties, and ease of parameterization (Jankowski et al., 2023; Carvalho et al., 2025). In this work, we focus on spline-parametrization of control trajectories. Spline-based trajectories ensure smooth and feasible control inputs that satisfy constraints and objectives inherent to MPC frameworks.
@@
[…]
@@
where $B_{i,p}:{\mathbb{R}}\to{\mathbb{R}}$ maps the time $t$ to the weighting of the $i$th control point, depicting the $i$th control point weight for blending (i.e., as with a probability distribution over $i$). Hence, the control waypoint ${\bm{u}}(t)$ is always in the convex hull of control points. The B-spline functions $B_{i,p}(t)$ are fully specified by a non-decreasing series of time knots $0=t_{1}\leq\ldots\leq t_{M}=T$ and the integer polynomial degree $p\in\{0,1,\ldots\}$ by
@@
$B_{i,0}$ are binary indicators of $t\in[t_{i},t_{i+1}]$ with $1\leq i\leq M-1$. The 1st-degree B-spline functions $B_{i,1}$ have support in $t\in[t_{i},t_{i+2}]$ with $1\leq i\leq M-2$, such that $\sum_{i=1}^{M-2}B_{i,1}(t)=1$ holds. In general, degree $p$ B-spline functions $B_{i,p}$ have support in $t\in[t_{i},t_{i+p+1}]$ with $1\leq i\leq M-p-1$. We need $K=M-p-1$ control points ${\bm{z}}_{1:K}$, which ensures the normalization property $\sum_{i=1}^{K}B_{i,p}(t)=1$ for every degree.
@@
Motivation. Spline representation provides several benefits. (i) It ensures smooth control trajectories (Watson & Peters, 2023) with guaranteed continuity of positions, velocities, and accelerations under mild assumptions of the dynamics. (ii) Spline simplifies complex trajectories through a few control points and efficiently incorporates constraints and boundary conditions, enabling efficient learning and planning by dimension reduction (Carvalho et al., 2025). (iii) It enables easy numerical optimization thanks to differentiable and convex representations.
@@ 3.1 Tensor Sampling
Figure 2: Illustration of different tensor interpolations on an evenly spaced graph with $M=3,N=9$. With a higher B-spline degree, the control trajectories exhibit more smoothness and conservative behavior, while Akima-spline aggressively and smoothly tracks control-waypoints. Note that we do not consider boundary conditions for B-spline interpolation in tensor sampling.
Inspired by (Le et al., 2025), we utilize the random multipartite graph as a tensor discretization structure to approximate global path sampling.
@@
Consider a graph ${\mathcal{G}}(M,N)=(V,W)$ on control space ${\mathbb{U}}$, the node set $V=\{L_{i}\}_{i=1}^{M}$ is a set of $M$ layers. Each layer $L_{i}=\{{\bm{u}}_{j}\in{\mathbb{U}}\mid{\bm{u}}_{j}\sim\textrm{Uniform}({\mathbb{U}})\}_{j=1}^{N}$ contains $N$ control-waypoints sampled from a uniform distribution over control space (i.e., bounded by control limits). The edge set $W$ is defined by the union of (forward) pair-wise connections between layers
@@
Sampling from ${\mathcal{G}}(M,N)$. The graph nodes are represented as the control-waypoint tensor for all layers ${\bm{Z}}\in{\mathbb{R}}^{M\times N\times n}$, within the control limits. To sample a batch of $B\in{\mathbb{N}}^{+}$ control paths with a horizon $T$, we subsample with replacement ${\bm{C}}\in{\mathbb{R}}^{B\times M\times n}$ from the set of all combinatorial paths in ${\mathcal{G}}$ (cf. Algorithm˜1) and further interpolate ${\bm{C}}$ into control trajectories ${\bm{U}}\in{\mathbb{R}}^{B\times T\times n}$ with different smooth structures, e.g., using Eq.˜6 or Eq.˜8. Sampling with replacement is cheap ${\mathcal{O}}(MN)$, while sampling without replacement is ${\mathcal{O}}(N^{M})$. Sampling without replacement adds overheads due to re-indexing or tree-traversing to get batch of sequence indices (depending on the low-level implementation of sampling) but offers better diversity. In practice, we use sampling with replacement, which does not really affect diversity (see last row in Fig.˜1), is faster and scales well with JAX vectorized operations. To see this, we have $N^{M}$ combinatorial paths in ${\mathcal{G}}(M,N)$, each path in ${\mathcal{G}}(M,N)$ has uniform $1/N^{M}$ mass, and the probability of sampling the same paths is small.
@@
[…]
@@
Control Path Interpolation. Straight-line interpolation can be realized straightforwardly by simply probing an $H=\lfloor T/M\rfloor$ number of equidistant points between layers, forming the linear coverage trajectories ${\bm{U}}\in{\mathbb{R}}^{B\times T\times n}$. However, there exist discontinuities at control waypoints using linear interpolation. Hence, we motivate spline tensor interpolations for sampling smooth control paths (cf. Fig.˜2).
@@
Given two time sequences $t_{i}=i/(M+p+1),\,i\in\{0,\ldots,M+p\}$ and $t_{j}=j/T,\,j\in\{0,\ldots,T-1\}$, the B-Spline matrix ${\bm{B}}_{p}\in{\mathbb{R}}^{M\times T}$ can be constructed recursive following Section˜2.2, with index $i,j$ corresponding to the element ${\bm{B}}_{i,p}(t_{j})$. Then, the control trajectories can be interpolated by performing einsum on the $M$ dimension of ${\bm{B}}_{p}\in{\mathbb{R}}^{M\times T},{\bm{C}}\in{\mathbb{R}}^{B\times M\times n}$, resulting ${\bm{U}}\in{\mathbb{R}}^{B\times T\times n}$.
@@
Definition 3 (Akima-spline Control Trajectories (Le et al., 2025)).
Given the time sequences $t_{i}=i/M,\,i\in\{0,\ldots,M-1\}$ representing the $M$ layer time slices and the control points ${\bm{C}}\in{\mathbb{R}}^{B\times M\times n}$, the Akima polynomial parameters ${\bm{A}}\in{\mathbb{R}}^{B\times(M-1)\times 4\times n}$ can be computed following Eq.˜10 in batch. Then, given the time sequence $t_{j}=j/T,\,j\in\{0,\ldots,T-1\}$, the control trajectories ${\bm{U}}\in{\mathbb{R}}^{B\times T\times n}$ are interpolated following polynomial interpolation Eq.˜8.
@@ 3.2 Path Coverage Guarantee
We analyze the path coverage property of ${\mathcal{G}}(M,N)$ for sampling control paths with linear interpolation. In particular, we investigate that any feasible path in the control space can be approximated arbitrarily well by a path in the random multipartite graph ${\mathcal{G}}(M,N)$ as the number of layers $M$ and the number of waypoints per layer $N$ approach infinity.
@@
Let $u\in{\mathbb{F}}$ be any control path and ${\mathcal{G}}(M,N)$ be a random multipartite graph with $M$ layers and $N$ uniform samples per layer (cf. Definition˜1). Assuming a time sequence (i.e., knots) $0=t_{1}<t_{2}<\ldots<t_{M}=1$ with equal intervals, associating with layers $L_{1},\ldots,L_{M}\in{\mathcal{G}}(M,N)$ respectively, then
@@
In intuition, Theorem˜1 states the support of path distribution ${\mathcal{G}}(M,N)$ approximates ${\mathbb{F}}$ and converges to ${\mathbb{F}}$ as $M,N\to\infty$. Thus, sampling paths from ${\mathcal{G}}(M,N)$ provides a tensorized mechanism to efficiently sample any possible path from ${\mathbb{F}}$, which allows vectorized sampling.
@@
To quantify the exploration level of tensor sampling, one standard way is to investigate its path distribution entropy. In intuition, when $M,N\to\infty$, tensor sampling entropy also approaches infinity due to uniform sampling per layer, which is further discussed in Section˜A.2. Note that this theory investigation serves as a guiding principle, while practical settings trade off entropy with success rate and runtime feasibility.
Practical Settings. We typically only set $M<T$. In principle, increasing $N$ should enable finer-grained exploration over the trajectory space. However, we observe diminishing returns when $N$ increases while keeping the total number of sampled trajectories $B$ fixed (cf. Fig.˜8). Intuitively, the underlying graph becomes denser, and the number of explored paths remains constant, resulting in only marginal performance improvements. Therefore, we recommend choosing $N$ proportionally to $B$ and within the bounds of available GPU memory, to maintain computational efficiency without oversampling from a small sample size.
@@ 3.3 Algorithm
Here, we present the overall algorithm combining tensor and local sampling with smooth structure options in Algorithm˜2. We propose a simple mixing mechanism with $\beta\in[0,1]$ by concatenating explorative and exploitative samples, forming a control trajectory tensor ${\bm{U}}$ (Line 4-8). We include the current nominal control for system stability at the fixed-point states (e.g., for tracking tasks) (Howell et al., 2022). Using simulators that allow for parallel runs (Todorov et al., 2012; Makoviychuk et al., 2021), rollout and cost evaluation can be efficiently vectorized (Line 9).
To tame the noise induced by tensor sampling, we modify the CEM update with softmax weighting on the elite set, for computing the new weighted control means and covariances similar to the MPPI update rule (Williams et al., 2017). We observe that this greatly smoothens the update over timesteps (Line 11-13) (cf. Fig.˜9). Finally, similar to Howell et al. (2022), we send the first control of the best candidate, since this control trajectory is evaluated in the simulator rather than the updated mean ${\bm{\mu}}$. Notice that we have fixed tensor shapes based on hyperparameters, for all subroutines of sampling, rolling out, and control distribution updates. Algorithm˜2 can be JAX jit and vmap over a number of $R$ model perturbations $\{f_{j}({\bm{x}},{\bm{u}})\}_{j=1}^{R}$, for efficient online domain randomization, while maintaining real-time control (cf. Section˜A.5).
@@
1 Choose interpolation type Linear, or B-spline (Definition˜2), or Akima (Definition˜3).
@@
5   Sample control waypoints ${\bm{C}}\in{\mathbb{R}}^{P\times M\times n}$ with $P=\lfloor\beta B\rfloor$, using Algorithm˜1.
@@
[…]